OpenStudy (anonymous):

Solve the system of equations using substitution. y=x^2+3x+2 y=2x+3

OpenStudy (shamil98):

so use substitution. 2x +3 = x^2 + 3x + 2

OpenStudy (anonymous):

just plot that y in \[ 2x+3=x^2+3x+2\]

OpenStudy (anonymous):

beat me!

OpenStudy (anonymous):

^^

OpenStudy (shamil98):

get everything on one side, then use the quadratic formula.

OpenStudy (anonymous):

i got x^2+x-1 but i cant factor that.

OpenStudy (shamil98):

a = 1 b = 1 c = -1 quadratic formula \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

Thanks.